Hanjina:

Illustrate the terms variables constant coefficient exponent degree of the polynomial for this expression by labelling the parts of the algebraic expressions: 1.25x² +6 5x³-20x² + x - 4

Timmyspu:

Welcome to Question Cove @hanjina. So what do we do first?

Timmyspu:

Yes I can so what is the first thing we have to do in the polynomial expression?

RinaRinaBoBina:

I believe you are right! So, you might wanna use pemdas to make this easier. (Parentheses, Exponents, Multiply, Divide, Add, Subtract!)

RinaRinaBoBina:

ok, so you don't have any parentheses, but you do have exponents, so take care of all the exponents first

RinaRinaBoBina:

I'm assuming this:6 5x³| is a 6.5x^3 ?

RinaRinaBoBina:

Anyway, you should have gotten: 1.25^2 = 1.5625x 6.5^3= 274.62500x 20^2=400x

RinaRinaBoBina:

So now instead of 1.25x² +6 5x³-20x² + x - 4, you have: 1.6x+274.625x-400x+x-4 If you deal with the x's first, your equation should turn into -122.77500-4 4 doesn't have an x, so you can't do anything with the two remaining numbers. If this is wrong I am so sorry. I literally just did this question in class yesterday, and the teacher said I got it right, but hey everyone makes mistakes. Good luck!